Shift report 20000630 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Jun 30 2000 - 15:57:16 EDT

  • Next message: http account: "Shift report 20000630 16:00-24:00"

    ------------------------------------------------------------
    Date: Fri Jun 30 15:57:16 2000, Shift: 08:00-16:00
    Supervisor: fv, with: KH,FJ,EJK,RD -- parttime
    Log book#2 pages 99 - 100
    
    Shift Summary:
    ------------------------------------------------------------
    Machine development- few attempt to accelerate many bunches, but no luck.
    Two Booster Main Power failures.
    The plans is soon to go back (if poosible) a more stable running for
    the weekend. we will see.
    
    Acheivements:
    The Ntuple summary plots for tof hits and Nses' has been change to
    be for trigger 6 (good BB-collisions events).
    Introduced students to operating experimen.
    
    Problems:
    All HV was still complete off following the Fault studies
    yeaterday. Found problem with BBL 4032 MF. Power was recysled
    (see page98) and works peoperly by now.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Jun 30 2000 - 16:01:31 EDT